APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday December 30, 2005 (MJD 53735) ---=== OBSERVING TEAM ===--- Swing: Dan Long Night: Viktor Malanushenko Elena Malanushenko John Barentine ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Battling mediocre seeing and soupy skies all night long, we managed to finish two plates (2343 and 2296) and obtained signal on four more (1543, 2067, 2051 and 2004). There was one problem involving a missing science exposure on plate 2051. See Log for details. ---=== OBSERVING LOG ===--- Afternoon: No problems during the afternoon checkout. Spectrograph focus checked but not changed. Still set for 3 to 4 C. Details in problem section. Looking into a couple items from last night's log: Checked enclosure and found it was in the normal operational state. Checked the rotator, by making several slews between 0 and 90, some at stow, some at 90, some moving between 0 and 90. There was the usual "failed to identify fiducial" at the start since it hadn't seen two yet, and I saw a couple failures to identify during a slew, but nothing excessive. The only thing I saw which might be a little out of the ordinary was an occasional faint deep rumble when the rotator was moving and higher motor current than I remember. It was typically reading around 7000, varying about +/- 3000. The highest numbers were just over 10,000 and seemed to occur only between 40 and 45 degrees. If I remember correctly that the current limit is 12,000 there seems a possibility of the rotator going out of closed loop if it goes a little higher when it gets colder etc. 23:45Z Doors Open, Louvers Open, Fans On 00:20Z Enclosure Open Night: The clouds thinned somewhat during twilight and by the time of the gotoField on the first cartridge of the night, the sky was mostly clear overhead. 00:40Z Cart 1 Plate 1543 (tile 9449, lowz74) We begin with a total (S/N)^2 of 5.2 on this plate from last night's observations. GSOGTF. While waiting for twilight to end, we had an error message in watcher involving the rotator: rotator (addoffset: Max accl. for rotator exceeded: 275/1000 (2005-12-31 00:50:50Z sdssmcp): 2005-12-31 00:50:58Z ) This occurred while we were tracking (i.e., not during a slew). Our first science exposure on this plate was for 900 s, giving a low (S/N)^2 of about 3. This matched well our results from the previous night, suggesting it was not clouds but faint objects on this plate that kept us from getting a better result. For the second and subsequent exposures, we used 1200 s. We obtained exposures totalling 5700 s on this plate before it expired at 03:14Z, but did not quite have enough signal to call this lowz-plate done. The clouds stole just enough signal to assure that. Sos asked for, asn we took, a postCalib before moving on. 03:20Z Cart 8 Plate 2067 (tile 8033, seguefaint105) From last night's observations we have a low total (S/N)^2 of 19.2 already on this plate. There was no repeat of last night's problem with instChange involving the rotator -- yee haw! GSOGTF. Clouds thickened during our first exposure, for 1200 s, and the result was a low (S/N)^2 of only 2.1, dangerously close to invalidating the exposure. We upped the time on subsequent exposures to 1500 s even though we still weren't really getting much. Toward 05:00Z we did a final 1200 s exposure that ran 10 minutes past the plate expiration, but 1) there was a 24-minute gap between this plate's expiration and the availability of the next plate and 2) efficiency was down due to thickening clouds. That exposure (36065) was declared bad because of (S/N)^2 < 2. We moved on. 05:30Z Cart 7 Plate 2051 (tile 8017, segue104) This plate already had a low total (S/N)^2 of 11 from previous observations. GSOGTF. A first 1200 s exposure gave a low (S/N)^2 of 2.7 in b1. Viktor issued a doNextScience command from his SOP session (as evidenced by SOPGUI) at 05:58Z, then a stopGuider command was recorded at 06:24Z. However, no such exposure (expected sequence number 36071) ever appeared on disk. sos_redo claimed there were no files to re-reduce. There is also no evidence in the Activity section of the log that a spectro exposure was underway at this time; in fact, there is an "end spectroScience" entry at 06:01Z, then nothing else until instChange at 06:24. For whatever reason, we lost about 25 minutes on this phantom exposure. The previous science exposure (36070) did not give us enough signal to declare this plate done. 06:35Z Cart 5 Plate 2296 (tile 1664, chunk123) No GSOGTF, but we found an fk5InFiber number 10. Clouds had thinned out by this plate but seeing was worse than earlier, and efficiency down to 70%. We did 2 x 1200 s exposures and backed off the third to 900 s. Finished the plate with a low (S/N)^2 of exactly 15.0. 08:06Z Cart 4 Plate 2343 (tile 1669, chunk123) GSOGTF. Efficiency plummeted during the first exposure as clouds thickened considerably. That exposure, for 900 s, did not have sufficient signal so we declared it bad and tried an 1800 s exposure next. That one also failed to meet the minimum (S/N)^2 requirement. We decided to continue on since it looked as though it might be clearing a little. Efficiency was up for the third exposure to between 50 and 70 percent. Waiting out the clouds seemed to work as our third exposure was valid. We did a total of 4 x 1500 s exposures which were good, finally finishing the plate. We also did the SoS-recommended postCalib. r2 is the lowest of the lot at a combined (S/N)^2 of 12.9, but the others are above 15 (b1 and b2 are > 30), so we are calling this one done. 11:15Z Cart 3 Plate 2004 (tile 1460, chunk100) GSOGTF. Sky looked better at the start of this plate than it had throughout the previous plate. We did two exposures here, but had to stop before the onset of twilight because of the zenith avoidance of this plate. We watched the telescope altitude very carefully and shortened our second exposure from 1500 s to 900 s in order to finish before we got much above 85 degrees. This plate's RaCen=32.16, so it spends almost 40 minutes inside the zone of avoidance, transiting at about 12:35Z. That left insufficient time coming down away from the zenith to start an exposure that would finish before 15 minutes past twilight. The low (S/N)^2 total for our two exposures is 10.9; we should finish this easily with one exposure tomorrow night. Closed for the night at 12:25Z. endNight rejoiced ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 5963 21:00Z 21:09Z 100 O 123.71 125.97 29 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): yes QA Procedures Done (y/n): yes UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 36045, plate -9999 ------- 21:03 36045 0.0 bias ----- sequence 36046, plate 1543 ------- 21:13 36046 10.0 flat 21:15 36047 2.0 arc 21:20 36048 2.1 arc 21:22 36049 2.1 arc ----- sequence 36050, plate -9999 ------- 23:31 36050 0.0 bias ----- sequence 36051, plate 1543 ------- 23.7 17.9 29.1 22.0 01:12 36051 10.0 flat 01:15 36052 2.0 arc 01:36 36053 900.1 target 01:59 36054 1200.1 target 02:22 36055 1200.1 target 02:45 36056 1200.1 target 03:08 36057 1200.1 target 03:11 36058 10.0 flat 03:14 36059 2.0 arc ----- sequence 36060, plate 2067 ------- 20.4 16.3 19.8 14.6 03:29 36060 10.0 flat 03:31 36061 2.0 arc 03:54 36062 1200.1 target 04:22 36063 1500.1 target 04:50 36064 1500.1 target 05:15 36065 1200.2 target 05:18 36066 10.0 flat 05:21 36067 2.0 arc ----- sequence 36068, plate 2051 ------- 2.7 1.5 6.0 6.4 05:34 36068 10.0 flat 05:37 36069 2.0 arc 06:00 36070 1200.1 target ----- sequence 36071, plate 2296 ------- 17.9 15.0 17.9 15.7 DONE 06:52 36071 10.0 flat 06:54 36072 2.0 arc 07:17 36073 1200.1 target 07:40 36074 1200.1 target 07:58 36075 900.1 target ----- sequence 36076, plate 2343 ------- 34.6 15.7 30.7 12.9 DONE 08:11 36076 10.0 flat 08:14 36077 2.0 arc 08:42 36078 900.1 target 09:10 36079 1500.1 target 09:38 36080 1500.1 target 10:06 36081 1500.1 target 10:34 36082 1500.1 target 11:01 36083 1500.1 target 11:06 36084 10.0 flat 11:08 36085 2.0 arc ----- sequence 36086, plate 2004 ------- 19.7 10.9 16.8 11.2 11:23 36086 10.0 flat 11:25 36087 2.0 arc 11:54 36088 1500.1 target 12:11 36089 900.1 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 01:18Z 1 1543 -11.46 0.0013 56.77 0.0020 170.27 -0.0149 0.999970 03:36Z 8 2067 -15.36 0.0004 47.61 0.0012 166.90 -0.0172 1.000070 05:42Z 7 2051 -30.13 -0.0009 46.90 0.0008 154.81 -0.0313 1.000030 06:56Z 5 2296 90.71 -0.0000 47.84 0.0030 246.97 0.0000 1.000230 08:26Z 4 2343 79.67 0.0010 60.99 0.0046 244.43 0.0020 1.000210 11:30Z 3 2004 91.71 0.0008 76.19 0.0040 262.81 0.0337 1.000180 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 01:18Z 1 1543 0.99997 296 1486 80 -12 56.8 3.4 9 270 -- 1.2 03:36Z 8 2067 1.00007 -691 746 130 -15 47.6 1.6 9 296 -- 1.5 05:42Z 7 2051 1.00003 -296 1026 80 -30 46.9 3.3 13 260 -- 1.8 06:56Z 5 2296 1.00023 -2271 -632 60 91 47.9 2.0 13 261 -- 1.8 08:26Z 4 2343 1.00021 -2074 -566 0 80 60.9 2.2 12 262 -- 2.0 11:30Z 3 2004 1.00018 -1778 -289 60 92 76.3 1.5 15 264 -- 1.8 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 21:01Z 44 5 9 281 (W) 445 - Patchy Cirrus 21:32Z 41 1 11 299 (WNW) 296 - Patchy Cirrus 22:03Z 42 0 15 305 (NW) 231 - Patchy Cirrus 22:37Z 44 -1 11 292 (WNW) 175 - Patchy Cirrus 23:09Z 40 -3 12 331 (NNW) 161 - Cirrus 23:41Z 39 -5 7 337 (NNW) 145 - Cirrus 00:13Z 39 -6 16 305 (NW) 113 - Cirrus 00:54Z 38 -8 8 272 (W) 111 - MstlyCldy 01:24Z 38 -8 7 270 (W) 118 - MstlyCldy 01:54Z 35 0 6 264 (W) 207 - MstlyCldy 02:25Z 35 7 8 263 (W) 410 - PtlyCldy 02:55Z 34 9 7 266 (W) 588 - PtlyCldy 03:26Z 34 10 7 281 (W) 613 - PtlyCldy 03:56Z 36 9 9 296 (WNW) 713 - PtlyCldy 04:27Z 36 6 8 273 (W) 521 - PtlyCldy 04:57Z 36 3 8 264 (W) 448 - PtlyCldy 05:27Z 37 0 14 263 (W) 308 - PtlyCldy 05:58Z 36 1 11 263 (W) 325 - PtlyCldy 06:29Z 35 1 12 257 (WSW) 330 - PtlyCldy 06:59Z 35 2 9 261 (W) 369 - PtlyCldy 07:30Z 35 2 10 252 (WSW) 367 - PtlyCldy 08:02Z 36 4 10 263 (W) 361 - PtlyCldy 08:32Z 36 3 12 264 (W) 401 - PtlyCldy 09:03Z 35 2 13 264 (W) 338 - PtlyCldy 09:35Z 36 2 13 263 (W) 410 - PtlyCldy 10:06Z 35 1 18 267 (W) 284 - PtlyCldy 10:36Z 35 2 15 263 (W) 379 - PtlyCldy 11:08Z 34 4 12 259 (W) 378 - PtlyCldy 11:41Z 34 5 16 266 (W) 491 - PtlyCldy 12:13Z 34 6 15 285 (WNW) 499 - PtlyCldy 12:45Z 34 5 19 284 (WNW) 516 - ---=== TELESCOPE STATUS ===--- 23:45Z Doors Open, Louvers Open, Fans On 00:20Z Enclosure Open 12:25Z Enclosure on, doors and louvers closed, fans off Telescope parked at 121,30 with Cart 1 (Plate 1543) mounted Counterweights at 245 Spectro autofill reconnected No interlocks bypassed Dewar weights at 12:45Z Imager: 139lbs 19psi Spectro: 86lbs 27psi (Note: Spectro dewars were force-filled at about 12:30Z. The big dewar should last until the afternoon.) ---=== SOFTWARE USED ===--- IOP/SOP: v4_9_2a Watcher: v2_31_0 MCP: v5_26_0 TPM: tpm_v3_1_0 AstroDa: v15_10 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_09 SoS: v5_1_0 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8350 -5.8440 Axial B -5.6770 -5.6770 Axial C 0.9390 0.9420 Trans D -10.0000 -9.9940 Lateral E 10.3632 10.3759 Lateral F 11.9507 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4711. -2457. 2997. -7825. 3296. 3024. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: 5.0 Alt.: 30.059746 MIGS TONIGHT NOMINAL Axial A 1.5480 1.5420 Axial B 1.0650 1.0470 Axial C 1.1840 1.1650 Trans D -0.1290 -0.1240 GALILS Commanded: 1592417. 1602036. 1553249. -6800. -5900. Actual: 1593447. 1603642. 1554316. -6820. -5882. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: 1256.02 15.75 -0.15 -49.66 136.81 ---=== PROBLEMS IN DETAIL ===--- Spectro Focus Details: sp1 sp2 Coll_motor_A 6000 -4490 Coll_motor_B 6000 -7770 Coll_motor_C 6000 -7100 im> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 2.915 tpm_TM_TelairAvg.VALA 4.350 b1 r1 b2 r2 Mean Offset -.01 -.02 -.41 .32 Recommended Piston 228 -2956 Indicated best focus temperature between 3 and 4 C Note: Best focus temperature can only reliably be determined by comparing against PSS temp using sp1. When the telescope has been at stow for an extended period (say all day), sp2 gets quite a bit colder (as much as 2 or 3 C) presumably due to the boil off of the camera dewars above the spectrograph body in this orientation. Note: When collimate is run in version v5_1_0 there is a warning message "speclog_version: Command not found." It doesn't appear to effect the calculations but is not seen in v4_10_7.